OpenStudy (anonymous):

Dale has 6 yard of rope. He cuts 16 inches off. What is the length of the remaining rope?

OpenStudy (anonymous):

1yard = 36inches, take 36*6 and subtract 16 from that. Hope that helps!

undeadknight26 (undeadknight26):

what he said...if you want to put that back in yards then: 36 x 6 = 216 - 16 = 200 200/36 = 5.55